OpenStudy (anonymous):

Suppose a matrix B' is obtained from a matrix B my multiplying the j-th row of B by a non-zero scalar k. Prove that Row(B) = Row(B')

OpenStudy (amistre64):

the rows represent coeffs of equations 2x + 3y = 5 for example creates a line on the graph n(2x + 3y = 5) 2nx + 3ny = 5n represents the same equation scaled by some number if we use say, elimination, techniques we should get the result that 0=0 2x + 3y = 5 ; *-n 2nx + 3ny = 5n -2nx -3ny =-5n 2nx+3ny = 5n --------------- 0 + 0 = 0 in other words, they are the same line and still retain all the same points.

OpenStudy (amistre64):

scaling an equation alters its "form", not its intrinsic values